Need problem solving process and the stepsversion 2 Questions 11 13 Suppose that when you buy a smartphone, you can also buy a protection plan at $21. The protection p
Answer #1

Profit profit if no compensation = 21 if compensation = 21-100 =8-19 - 2 C) Expected pagit = 20 + (-74) + 80 (21) 2017 + (20)

Ans a $1 1 ) E (P2) = 20 (-70)² + 30 (21)2 = 1248.2 + 352.8. Variance - Var (P = E(P²) - (66f)) = 1601 -

DI Pg. Standard deviation = 1600 = 40 z statistic o - 1 - 1x10 - 0.25 . It will pllow normal distribution as nslod o bolo is

makes pugit , so above o is profit only Sering Dalue Probability of company making profit : 0.5817 P ( 2 z 3-0.25) = D B tabl

Bo Following same process as lost question 2 statistic = = cat) (20)

2 statistic = Probability company manas profit P ( 2 ² -0.5 = 0.6915 This becomes 69.15)